if period of function cos(nx).sin(5x/n) is 3 pi, then number of integral values of n are ?

Grade:12th pass

8 Answers

jitender lakhanpal
10 years ago

hi subhadeep

the function can be written in the form sin (A+B) + sin (A-B) using trigonometric identity

then this is of the form of sum of two functions then find the L.C.M of the two arguments

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

 

jitender ''

askiitians expert

SUBHADEEP ROY
10 years ago

your approach is correct but how to solve further......taking the LCM and finding values of n ???

that is my question........even i already knew the approach of sin (a+b)+sin(a-b)........if possible tell me the method after that .............i am stuck in taking the LCM.........and what is the aswer u r getting???

jitender lakhanpal
10 years ago

hi subhadeep

the answer is no integral solution

actually we dont have to use the product to sum identity in trigonometry

the product of any two periodic function is also periodic

we just have to find L.C.M of their fundamental period

fundamnetal period of cos(nx) is T1 = 2 pie / n -------1

and other function the fundamental period is T2 = 2npie/5------------2

their ratio (T1/T2) = 5 / n^2

 n^2 T1 = 5 T2 = T = 3 pie ------ given   (T is the composite fundamental period )

T1 = 3 pie / n^2       T2 = 3 pie / 5

putting the values of T1 and T2 we get n = 3/2

so no integral solution

SUBHADEEP ROY
10 years ago

why did u take the ratio of T1 and T2 and then cross multiplying equated to T = 3 pi ..........??

the logic ???

jitender lakhanpal
10 years ago

hi subhadeep

if c is L.C.M of a , b

a, b can be written as

c = ma       , c = nb 

where m,n are integes

so c = ma = nb

that is why i took the ratio of T1 , T2  and wrote in terms of T

and also it was necessary as we had to find the value of n

for any other help and insights you may join ASKIITIANS

 

jitender

SUBHADEEP ROY
10 years ago

but the answer given is 8 integral solutions ............

jitender lakhanpal
10 years ago

pls provide those integral values if they are  given in the answer

Y RAJYALAKSHMI
9 years ago
the integral values of n are ±1, ±3, ±5, ±15